OpenStudy (anonymous):

Assume the government cuts taxes by $5 million. If the MPC = 0.75, what is the maximum potential impact on real GDP according tot he Keynesian model?

OpenStudy (kira_yamato):

It is a infinite geometric progression. So k = -MPC/1-MPC (sum from one to infinity) = -0.75/0.25 = -3 Maximum impact = -3(-5) = 15 mil

OpenStudy (anonymous):

thank you!